OpenStudy (anonymous):

The price of an oven changes from $450 to $396 a.13.6% increase b.13.6%decrease c.12% increase d.12% increase

OpenStudy (tkhunny):

What is the AMOUNT increase?

OpenStudy (anonymous):

umm??

rebeccaxhawaii (rebeccaxhawaii):

well it an obvious answer pretty much common sense the numbers are decreasing so that would make it only one choice

OpenStudy (anonymous):

OK its either c or d

OpenStudy (anonymous):

I am lost someone help

rebeccaxhawaii (rebeccaxhawaii):

read my last answer it shouldnt be confusing that question literally gives you the answer

OpenStudy (anonymous):

ok

OpenStudy (tkhunny):

Answer my question. What is the AMOUNT of the Increase or Decrease?

OpenStudy (anonymous):

I don't know the amount

OpenStudy (tkhunny):

Can you subtract and find it?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

54

OpenStudy (tkhunny):

450 - 396 = 54. Perfect. Where did we start? 450 or 396?

OpenStudy (anonymous):

Yes $450 to $396

OpenStudy (tkhunny):

Okay, What percent of where we started is the change? In this problem, What percent of 450 is 54?

OpenStudy (anonymous):

45%

OpenStudy (tkhunny):

\(\dfrac{54}{450} = 0.12 = 12%\) How did you get 45%

OpenStudy (anonymous):

ohh opps

OpenStudy (anonymous):

ok so its 12

OpenStudy (tkhunny):

It's a 12% what? Increase or Decrease?

OpenStudy (anonymous):

Decrease

OpenStudy (tkhunny):

The correct answer is 12% Decrease. This isn't on the list. Perhaps you mistyped answer choice D?

OpenStudy (anonymous):

yea i did

OpenStudy (anonymous):

thank you for helping me

OpenStudy (tkhunny):

Did you see what Rebecca was saying? If we look at it and see that it is a decrease, it would be quite simply to pick the only answer that actually says decrease! Unfortunately, that wasn't the real solution. Do you know where 13.6% comes from?

OpenStudy (tkhunny):

If you forget where you STARTED, you may accidentally calculate \(\dfrac{54}{396} = 13.6\)%. That would be wrong. Always use where you STARTED.